I have two more questions:
1. Two angles are supplementary, and the ratio of their measures is 7 to 2 what are the angle measures?
2. Two angles are complementary, and the ratio of their measures is 5 to 4. What are the angle measures?
: I have two more questions:
1. Two angles are supplementary, and the ratio of their measures is 7 to 2 what are the angle measures?
2. Two angles are complementary, and the ratio of their measures is 5 to 4. What are the angle measures?

Answer by nerdybill(1129)


1. Two angles are supplementary, and the ratio of their measures is 7 to 2 what are the angle measures?
When two angles are supplementary, the sum of the two angles equals 90.
7/9 * 90 = 70 deg
2/9 * 90 = 20 deg
2. Two angles are complementary, and the ratio of their measures is 5 to 4. What are the angle measures?
When two angles are complementary, the sum of the two angles equals 180.
5/9 * 180 = 100 deg
4/9 * 180 = 80 deg